# Thermoelectric effects of an interacting hadron gas in the presence of an external magnetic field

### v1

@@ IV Summary
The presence of temperature gradients from the central to the peripheral region of heavy-ion collision can lead to thermoelectric currents in the medium in the case of non-zero baryonic chemical potential. We have calculated several thermoelectric coefficients within the ideal and other interacting hadron resonance gas (HRG) models. For the first time for a hot and dense hadronic medium, we calculated the Thomson coefficient, which was introduced due to the temperature dependence of the Seebeck coefficient. For the completeness of our study, we have also calculated the thermoelectric figure of merit $ZT$ of the hadronic medium. We relate the gradients of baryon chemical potential to the gradients of temperature using the Gibbs-Duhem relation. During this modification, we observed the introduction of enthalpy in the Seebeck coefficient formula. Although the Seebeck coefficient $S$ for mesons vanishes due to equal and opposite charges, mesons still have a significant effect on the $S$ because of their contributions to enthalpy. We have also studied the effects of finite magnetic fields on the thermoelectric properties of the HRG medium. As the hadronic medium is a conducting medium, the presence of a magnetic field leads to Hall-like components due to the Lorentz force. In the absence of a magnetic field, the magneto-Seebeck coefficient $S_{B}$ merges into the Seebeck coefficient. We observed negative values $S_{B}$ for all cases, except for the high magnetic field ($eB=1.0~{}m_{\pi}^{2}$) and baryon chemical potential ($\mu_{B}$ = 0.40 GeV) at lower temperatures where $S_{B}$ is positive. On the other hand, the induced current in the direction perpendicular to the electric field is measured with the help of the normalized Nernst coefficient $NB$. The higher is the magnetic field value, higher is the value of $NB$. For all the cases, we observe that at higher temperatures, $NB$ approaches zero. The presence of a magnetic field also affects the thermoelectric figure of merit $ZT$. The higher is the magnetic field, higher is the nonmonotonic behaviour of $ZT$. The thermoelectric efficiency of the medium is the same for all the HRG models at the lower temperature with finite values of the magnetic field at a fixed $\mu_{B}$. However, it is highly deviated for each of the models at higher temperatures.
The current study focuses on estimating the thermoelectric coefficients of a hot and dense hadronic medium formed in heavy-ion collisions using different HRG models to study the impact of hadronic interactions. However, it will also be interesting to study the magneto-Thomson effect in hadron gas and in the QGP medium.
Acknowledgement
KS and KKP acknowledge the financial aid from UGC, Government of India. The authors gratefully acknowledge the DAE-DST, Govt. of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021- IITI (E-37123).

### v2

@@ IV Summary
The presence of temperature gradients from the central to the peripheral region of heavy-ion collision can lead to thermoelectric currents in the medium in the case of nonzero baryonic chemical potential. We have calculated several thermoelectric coefficients within the ideal and other interacting HRG models. For the first time for a hot and dense hadronic medium, we calculated the Thomson coefficient, which was introduced due to the temperature dependence of the Seebeck coefficient. For the completeness of our study, we have also calculated the thermoelectric figure of merit $ZT$ of the hadronic medium. We relate the gradients of baryon chemical potential to the gradients of temperature using the Gibbs-Duhem relation. During this modification, we observed the introduction of enthalpy in the Seebeck coefficient formula. Although the Seebeck coefficient $S$ for mesons vanishes due to equal and opposite charges, mesons still have a significant effect on the $S$ because of their contributions to enthalpy. We have also studied the effects of finite magnetic fields on the thermoelectric properties of the HRG medium. As the hadronic medium is a conducting medium, the presence of a magnetic field leads to Hall-like components due to the Lorentz force. In the absence of a magnetic field, the magneto-Seebeck coefficient $S_{B}$ merges into the Seebeck coefficient. We observed negative values $S_{B}$ for all cases, except for the high magnetic field ($eB=1.0~{}m_{\pi}^{2}$) and baryon chemical potential ($\mu_{B}$ = 0.40 GeV) at lower temperatures where $S_{B}$ is positive. On the other hand, the induced current in the direction perpendicular to the electric field is measured with the help of the normalized Nernst coefficient $NB$. The higher the magnetic field value, the higher the value of $NB$. For all the cases, we observe that at higher temperatures, $NB$ approaches zero. The presence of a magnetic field also affects the thermoelectric figure of merit $ZT$. The higher the magnetic field, the higher the nonmonotonic behavior of $ZT$. The thermoelectric efficiency of the medium is the same for all the HRG models at the lower temperature with finite values of the magnetic field at a fixed $\mu_{B}$. However, it is highly deviated for each of the models at higher temperatures.
The current study focuses on estimating the thermoelectric coefficients of a hot and dense hadronic medium formed in heavy-ion collisions using different HRG models to study the impact of hadronic interactions. However, it will also be interesting to study the magneto-Thomson effect in hadron gas and the QGP medium.
ACKNOWLEDGMENTS
K.S and K.K.P. acknowledge the financial aid from UGC, Government of India. The authors gratefully acknowledge the DAE-DST, Govternment of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021- IITI (E-37123).